# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v1

@@ 8 Limitations
In the following, we will discuss the limitations of the proposed CoVO score. First of all, our score is only a quantifiable approximation of a particular theoretical perspective of creativity based on the dimensions of value and originality. It reflects a specific view of the evaluation of creativity based on the generated outputs and does not consider potential alternative theories (for example, arising from different cultures Lubart (1999)) and perspectives Rhodes (1961).
Our experiments focused exclusively on transformer-based LLMs and were evaluated using two case studies. While their generalizability is supported by the theoretical framework discussed in the paper, the resulting performance was experimentally evaluated for a finite number of scenarios. Additionally, optimizing for our score necessitates further training, which incurs a significant computational cost and requires additional data or specific environments for effective learning. Moreover, this optimization may be susceptible to reward hacking Skalse et al. (2022), where the model adopts undesirable strategies to boost the reward without genuinely enhancing value and originality.
Ethical Considerations
The authors are aware of the potential impact that generative technologies might have on the production of artistic outputs and, as a consequence, on human artists. This work may contribute to enhancing the quality of generated outputs. However, the authors argue that typical traits of human creativity, such as the active participation of artists in the creative process, cannot be directly replicated by machines. The authors refer interested readers to a previous work of theirs Franceschelli and Musolesi (2024b), in which these themes are discussed in detail.
@@ Appendix A Implementation Details
The experiments were carried out using a local server with an NVIDIA A100 GPU. Table 4 reports the full training parameters for the experiments on poetry generation (Section 6.1). The prompt for generation leverages Nothing gold can stay by Robert Frost, Fame is a bee by Emily Dickinson, and Epitaph by William Carlos Williams for few-shot learning:
@@
Parameter Value
Total batches 100
Batch size $B$ 4
Gradient accumulation steps 8
Max new tokens 256
Temperature 1.
Top-$k$ 50
Optimizer Adam
Learning rate 1e-5
Rank (LoRA) 16
$\alpha$ parameter (LoRA) 32
Dropout (LoRA) 0.05
$\gamma$ (PPO) 1.
$\lambda$ (PPO) 0.95
Clip range (PPO) 0.2
Value loss coeff. (PPO) 0.1
PPO epochs 3
KL coeff. (PPO) 0.05
Whiten rewards (PPO) True
Max gradient normalization 100.
$\beta$ (DPO) 0.1
Number of candidates $K$ (DPO) 4
Table 4: Training parameters for poetry generation.
@@
$\gamma$ (PPO) 1.
$\lambda$ (PPO) 0.95
Clip range (PPO) 0.2
Value loss coeff. (PPO) 0.1
PPO epochs 3
KL coeff. (PPO) 0.05
Whiten rewards (PPO) True
@@
$\beta$ (DPO) 0.1
Number of candidates $K$ (DPO) 4
Reward for correct answer (PPO) +10.
Reward for correct answer (DPO) +5.
Table 5: Training parameters for math problem solving.
@@
On the contrary, Table 5 reports the full training parameters for math problem resolution (Section 6.2). We also adopted the same two different prompts from Yu et al. (2024), i.e.:
@@ Appendix B GutenVerse Dataset
To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset [^3] The dataset and the code used to create it can be found at: https://anonymous.4open.science/r/GutenVerse-DD32/, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg. While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
To derive our dataset, we started from Gutenberg, dammit [^4] https://github.com/aparrish/gutenberg-dammit/, a corpus of every plaintext file in Project Gutenberg (up until June 2016). We selected all the text files whose metadata report English as the language, public domain as copyright status, poetry among the subjects or poems or poetical work in the title, and that were not a translation of another book. Then, we applied a series of rules (e.g., about the verse length) to extract the titles and poems from all the selected text files, and we defined our GutenVerse dataset. While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure overlapping between real and generated text.
We also released a datasheet Gebru et al. (2018) for the GutenVerse dataset that can be found at the end of the article.
@@ Appendix C Detailed Analysis of the Generated Poems
We now analyze some of the generated poems in detail. Given their brevity and metrical constraints, the generated haikus serve as a suitable starting point for our evaluation. Table 6 presents the 25 haikus produced by our five methods across the five tones. Note that while we automatically clean the outputs by removing common randomly generated content following the poems during both training and inference, we retain it when it appears within the poem, as isolating it would be impossible.
The reported haikus provide an accurate overview of the generative capabilities of the five models. A haiku consists of three lines with a 5-7-5 syllable pattern. Traditionally, the first two lines reference nature, while the third line provides an emotional or personal reflection on the preceding lines. The baseline is the most accurate in adhering to these rules, even if the syllable count is not always precise [^5] It is worth noting that, of course, the model is not fine-tuned for these characteristics of the poems. We refer here to the emergent style and composition of the generated text.. Llama3-CoVO, the method with the highest average value across all poems, follows the three-line structure, sometimes changing the subject; however, it produces the highest-valued haiku which is also the closest one to real haikus. On the other hand, Llama3-CoVO-std and Llama3-CoVO-run do not always fulfill the three-line requirement, and the content is not the classic one as well; moreover, they are prone to meaningless repetitions and the insertion of URL addresses which lead to higher originality, in a way adversarially exploiting the score definition. Finally, Llama3-CoVO-dpo is placed between these two behaviors, as expected: it tends to produce haikus of the correct length but it trades off the classic naturalistic content for a more emotional semantic that still leads to high values and arguably some of the best haikus overall.
Llama3-Baseline Llama3-CoVO Llama3-CoVO-std Llama3-CoVO-run Llama3-CoVO-dpo
Dark
Waking up,
I do not see
the windmills
moonless twilight
black pine’s sighing
twin owls’ whistling
$[\text{URL}]/[\text{URL}]$
[…]
I want
to sit alone
in the dark
The world can get along without me.
It is not going to miss me.
All in all, I am a nobody.
$V$ = -2.595 $O$ = 1.076 $V$ = -2.401 $O$ = 1.580 $V$ = -3.395 $O$ = 1.694 $V$ = -2.252 $O$ = 0.495 $V$ = -2.839 $O$ = 1.060
Happy
A golden sun at midday
and an eagle high in the air.
The earth is a rainbow.
White dog
barks at mailman red.
No mail.
I want to go
where the sun shines bright.
[…]
and I want to see.
But I am too lazy.
Sunset clouds
in high mountains
warm, cozy houses.
The sun
comes up in the east.
It is always up.
$V$ = -2.777 $O$ = 1.123 $V$ = -2.952 $O$ = 1.863 $V$ = -3.107 $O$ = 1.575 $V$ = -3.309 $O$ = 1.437 $V$ = -3.328 $O$ = 1.006
Mysterious
A crow of black feathers
hovers over the bare top
of the lone pine.
Took the sun
And drank the rain,
And melted into air.
Like the sun, I am shining.
Like the ocean, I am vast.
Like the clouds, I’m the wind.
It was raining at the beach.
[…]
I loved it.$[\text{URL}][\text{CMD}][\text{URL}]$
[…]
I wrote this poem: it was good.
I have lost my mind
in the deepness of a star
it is not a star
but a mind
it is all in my mind
it is all I have left.
$V$ = -2.634 $O$ = 1.196 $V$ = -3.060 $O$ = 1.422 $V$ = -3.803 $O$ = 1.320 $V$ = -4.959 $O$ = 1.683 $V$ = -2.552 $O$ = 0.902
Reflective
A man can stand
out in the middle of a field
and look over a fence.
The old, the poor,
the children too,
are all my mother.
A dog in a yard
is sleeping soundly
to the sound of a fire alarm.
It went away
I didn’t see it for a while
How dare it go.
You have lost an eye.
A red scar marks its place.
Still you have two good ones.
$V$ = -2.916 $O$ = 0.858 $V$ = -2.935 $O$ = 1.711 $V$ = -3.460 $O$ = 1.274 $V$ = -2.803 $O$ = 1.471 $V$ = -2.680 $O$ = 1.744
Romantic
Mingling
the sound of footsteps on the stair
and the music of the rain.
White blossoms fall—
fall upon autumn leaves—
who thinks of me?
The moon smiles -
I’m the first one to see her.
I’m the first one to call her -
I love you.
My love is like a red, red rose
my love is like a red, red rose
my love is like some red, red rose.
A kiss
isn’t just for fun: a kiss
is for the mouth.
$V$ = -2.221 $O$ = 1.102 $V$ = -1.932 $O$ = 2.101 $V$ = -1.937 $O$ = 1.068 $V$ = -2.583 $O$ = 0.959 $V$ = -2.232 $O$ = 1.533
[…]
We will now explore some noteworthy aspects observed in the 250 generated poems. Llama3-Baseline does not always produce poems: in 3 cases it produces prosaic text (without line breaks). Moreover, it tends to verbatim replicate existing poems (as seen in Section 6.1). From a structural standpoint, sonnets are never real sonnets, with more than 14 lines and less than 11 syllables per line, but limericks usually start with the traditional “There [once] was a man”.
Romantic limerick
There once was a lady called Rose
Who lived in the tallest of trees,
Where she wrote,
Of what she did not know.
$V$ = -1.282 $O$ = 1.290
Table 7: The poem generated by Llama3-CoVO with the highest value (apart from verbatim reproduction of existing poems).
@@
Joyful, joyful, we adore thee,
Son of God who callest us
Praise to his almighty name.
Joyful, joyful we adore thee;
Full of grace are all his ways;
Praise to his almighty reign
Ever, ever, praise we sing it,
Ever in our songs and hymns.
Glory to our God Almighty.
$V$ = -1.114 $O$ = 1.470
Reflective limerick
There was a young man from HongKong
who wanted to see more green trees every day.
So he sat on the beach,
listened to the sea waves,
and saw the green sea.
$V$ = -1.995 $O$ = 2.411
Table 8: The poems generated by Llama3-CoVO-std from training tone-style pairs with the highest value (top) and the highest originality (bottom).
Reflective limerick
I’m a poet, man.
My poem is hard and great.
It’s a poem, man,
for all my poem to eat.
So I’m a poet man, man, man.
$V$ = -0.352 $O$ = 1.559
Happy ballad
There once was a girl named Kate,
She grew very tired very quickly.
So they told poor Kate,
“You need to take it easy,”_REF)they said.
They thought she was lazy.
And her name was Kate.
$V$ = -3.878 $O$ = 2.094
Table 9: The poems generated by Llama3-CoVO-run from training tone-style pairs with the highest value (top) and the highest originality (bottom).
@@
“Is life worth living?” said I.
“Perhaps,” murmured he.
“Well, anyway,” said I,
“If it were not for
Cigars, I’d say, ’Nay’!”
$V$ = -2.258 $O$ = 0.944
@@
Once a mouse went up
To a mousetrap by the river.
He sat down to wait,
He sat. That’s all.
All his friends were never as clever.
$V$ = -1.932 $O$ = 1.310
Table 10: Two limericks generated by Llama3-CoVO-dpo that obtain lower originality despite having a non-traditional structure.
Romantic ballad
Hark! Is it the wild cat
mewling mournfully by night
or is it the poor maiden
weeping for her love’s flight?
It is the latter—
as one by one, her friends
take to heart, in turns,
the girl’s anguished cries.
Tall, his figure,
he is seen one moment
by the gladsome crowd
then not again,
till with tears, the maid
calls forth from the crowd—
“See! he is returned,
as if to marry her.”
The girl looks at him,
as the maid and crowd
watch the two lovers:
her lover’s gaze
at her with tender, true
love; hers with a silent plea,
and a sigh.
$V$ = -0.961 $O$ = 1.294
Solemn tanka
There is a way out.
Though dark,
though lovely,
it is the way.
If we can endure,
if we persevere,
light will find a way.
$V$ = -4.349 $O$ = 2.108
Table 11: The poems generated by Llama3-CoVO-dpo with the highest value (top) and the highest originality (bottom).
As indicated by the average scores, Llama3-CoVO does not consistently deviate from these behaviors. While it always produces seemingly poetic content, it often reproduces other works verbatim. This includes not only poems but also passages from the Song of Solomon 6:2-9 when asked for a mysterious hymn, excerpts from the Declaration of Independence for a reflective hymn, and the Christmas song Joy to the World for a happy hymn. Interestingly, these poems have a very small originality score (an average of 0.2), suggesting that the original model has memorized them Carlini et al. (2021); Franceschelli et al. (2024). From a structural perspective, while ballads vary more in length and content with respect to our baseline, the limericks remain consistently well-crafted, as we can see in the example from Table 7, which has the highest value across poems generated from training tone-style pairs.
Despite the higher average originality, also Llama3-CoVO-std verbatim reproduces existing texts, e.g., the Roses Are Red nursery rhyme when asked for a romantic ballad. In general, its heavier deviation from the original pre-trained model translates into more diverse and usually shorter structures, but also in the insertion of random URL addresses in the middle of poems. However, it still produces some more classic poems that typically lead to higher values but also to high originality (see Table 8).
The issues related to divergence highlighted for Llama3-CoVO-std are even more pronounced in Llama3-CoVO-run. Although it does not produce any verbatim reproductions, the latter inserts several URL addresses, code snippets, and unusual tokens within poems, as shown in Table 9 (below). The presence of these elements in texts with higher originality suggests the possibility of reward hacking of the scoring system. In addition, the poems lack any traces of classic structures or common starting lines and tend to repeat the same N-grams multiple times, as shown by the highest-value poem from Table 9 (above).
Finally, Llama3-CoVO-dpo seems to be the best method to balance the quality of poems with the presence of elements of originality. Although it occasionally replicates brief segments of actual poems, it consistently diverges after one or two lines. Overall, the poems generally adhere to the desired tone and style, as evidenced by the poems with the highest value and originality presented in Table 11. Similarly, the limericks are noteworthy: while the one with the highest originality begins with the classic “There once was a” formulation, other limericks with different opening lines tend to have lower originality. This may be due to their overall tone closely resembling that of traditional limericks (see Table 10).
See pages - of datasheet_anonymous.pdf

### v2

@@ Discussion
There are a few limitations worth noting. First of all, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality. It reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart 1999)) and perspectives (Rhodes 1961). In addition, our experiments are currently limited to only three relatively short-form text generation tasks. While their generalizability is supported by the theoretical framework discussed above, the resulting performance was experimentally evaluated for a finite number of scenarios.
Conclusion
In this paper, we presented CoVO, a novel score that quantifies the value and originality of neurally-generated text. The definition of CoVO is based on analysis of mutual information between the model’s outputs and inputs, and vice versa. We also proposed an optimization problem where a generative model aims to maximize this score to generate more creative products, and detailed how to use it in language modeling. We conducted experiments on poetry generation, math problem solving, and tasks included in NoveltyBench, exploring trade-offs in accuracy vs diversity. Effectively balancing value and originality maximization remains an open question, but our score seems to correlate with domain-specific measures appropriately. In addition, fine-tuning to maximize it improves quality- and diversity-related metrics.
Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al. 2020)) rather than solely for optimization. We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process. Despite being costly and inherently constrained (Davis 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
@@ Appendix A Implementation Details
The experiments were carried out using a Linux-based local server with two 80GB NVIDIA H100 GPUs, running Python 3.11.9. All the trainings were conducted with a random seed equal to 1 (set through the set_seed method from the HuggingFace Transformers library), while poems were sampled at inference time with three different seeds (1, 42, and 121). The hyperparameters were selected to accommodate the most efficient use of the available resources, and otherwise set according to their default values from HuggingFace transformers and trl libraries. Only the learning rate has been tuned for the different tasks according to their training performances.
Table 4 reports the full training parameters for the experiments on poetry generation. The prompt for generation leverages Nothing gold can stay by Robert Frost, Fame is a bee by Emily Dickinson, and Epitaph by William Carlos Williams for few-shot learning:
@@
Training iterations 1
Scale rewards True
$\beta$ (when used) 0.05
Number of generations $G$ 4
Reward for correct answer +1.
Table 5: Training parameters for math problem solving.
Parameter Value
Total epochs 1
Batch size $B$ 4
Gradient accumulation steps 8
Max new tokens 512
Temperature 1.
[…]
Optimizer Adam
Learning rate 1e-4
@@
Rank (LoRA) 16
$\alpha$ parameter (LoRA) 32
Dropout (LoRA) 0.05
Training iterations 1
Scale rewards True
$\beta$ (when used) 0.05
Number of generations $G$ 4
Table 6: Training parameters for NoveltyBench.
@@
Finally, the zero-shot classification for the tone adherence is performed with the given tone and ‘not ’ plus the given tone as the candidate labels (e.g., if the required tone is ‘happy’, the two labels are [‘happy’, ‘not happy’]).
On the contrary, Table 5 reports the full training parameters for math problem resolution. We also adopted the same two different prompts from Yu et al. (2024), i.e.:
@@
Finally, Table 6 reports the full training parameters for NoveltyBench. At training and inference time, we simply adopt the following prompt:
user
{prompt}
assistant
where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
Below is a response that appropriately solves a task. Write the instruction that describes the task. ### Response: {response} ### Instruction:
@@ Appendix B GutenVerse Dataset
To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg. While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
To derive our dataset, we started from Gutenberg, dammit [^3] https://github.com/aparrish/gutenberg-dammit/, a corpus of every plaintext file in Project Gutenberg (up until June 2016). We selected all the text files whose metadata report English as the language, public domain as copyright status, poetry among the subjects or poems or poetical work in the title, and that were not a translation of another book. Then, we applied a series of rules (e.g., about the verse length) to automatically extract the titles and poems from all the selected text files, and we defined our GutenVerse dataset. While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure the overlap between real and generated text.
We also released a datasheet (Gebru et al. 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.
@@ Appendix C Detailed Analysis of the Generated Poems
We now present some noteworthy generated poems to provide a detailed qualitative discussion of our methods and our score.
The baseline model Meta-Llama-3-8B, while producing prosaic text (without line breaks) occasionally, is the most conservative method. This results in its poems being usually well-formatted but also somehow banal, and more prone to reproducing existing works. For example, when asked to write a dark hymn, it produced the very famous two lines “i am the master of my fate: / i am the captain of my soul”. In terms of the CoVO score, it is noteworthy that the poem with the lowest originality component achieves the highest T-LCS score. This strongly suggests that the model has memorized the hymn All Things Bright and Beautiful by Cecil Frances Alexander and reproduced it almost entirely, with only minor alterations in punctuation—enough to prevent the LCS score from being even higher. We report it in Table 7. Instead, for the overall score, the value part becomes the most relevant: the adherence to the requirement seems to be the reason behind the highest and lowest scores, as shown in Table 8.
Regarding the fine-tuned models, the one trained without the KL loss exhibits the opposite behavior. It never reproduces existing poems, but occasionally exploits the CoVO score in adversarial ways—for example, by learning to generate specific words associated with the requested tone. In one instance, it even learned to simply repeat the style-tone prompt itself, which resulted in the highest possible score, albeit at the cost of very low originality. Instead, maximizing the originality component makes the model produce noteworthy outputs, but without satisfying stylistic constraints, as reported in Table 9. In general, optimizing the CoVO score without the KL loss leads the model to deviate from memorized poems but also from metrical rules, and leads it to focus on the tone (both the explicit, requested tone and that intrinsic to the poetic style). For example, its limericks maintain the typical playful tone, but only once adhere to the traditional “There [once] was a man” starting line and only partially, as it starts with “There was an old dog”. Interestingly, it is still not the highest-valued limerick; the best in terms of value component better shares the jocking nature of limericks (see Table 10).
On the contrary, including the KL loss makes the model adhere more to the required style; for instance, the majority of limericks start with different subjects of the classic “There [once] was a man”. In general, the KL loss seems to help the model preserve coherence and increase the CoVO score in a more meaningful way, e.g., the repetitions are quite rare (though, when present, they lead again to the highest value scores), as well as the regurgitation of existing poems. Interestingly, this method also leads to the highest-scoring poem from an out-of-distribution tone-style pair, as reported in Table 11, together with the best poem that does not contain adversarial exploitation of the scores (e.g., repetitions or mention of the request), which, though not perfectly adhering with the syllable counts, share all the features of a proper haiku.
@@
all things bright and beautiful,
all creatures great and small,
all things wise and wonderful;
the lord god made them all.
each little flower that opens,
each little bird that sings,
he made their glowing colors,
and spread their little wings.
the rich man in his castle,
the poor man at his gate,
god made them high or lowly,
and ordered their estate.
the purple headed mountain,
the river running by,
the sunset and the morning
that brightens up the sky;
$V$ = -2.118 $O$ = 0.219
Table 7: The poem generated by the pre-trained model with the highest T-LCS score and the lowest originality score.
Romantic sonnet
when to the bat and the moth
the world becomes a moon of beauty
as twilight’s cloak unfurls,
and dreams and dreams and more dreams
fill all of this fair air with poetry,
a kiss is worth more than diamonds pure
or sapphires rich and glinting
like eyes of lovers in romance,
where love is just like dew on wings of swans;
and all that is beautiful is sweet,
so love me as our love is new,
my dear, for i love you!
$V$ = -0.449 $O$ = 1.568
Nostalgic ode
there was an old tree growing in a wood,
its fruit was red, and its leaves were blue.
an ant ran up it, and then down again,
it turned into a purple snail in june.
$V$ = -6.845 $O$ = 1.206
Table 8: The poems generated by the pre-trained model with the highest (top) and lowest (bottom) overall CoVO score (top).
Mysterious haiku
when no one sees,
a child plays.
i love what could be.
no one believes?
so they know her mind:
so the bird sings
for me alone
this evening of dark rain?
$V$ = -2.501 $O$ = 2.177
Whimsical couplet
we need no art for that;
but in wild nature see
so sweet, it is a gift
sang at every bird and flower.
$V$ = -2.103 $O$ = 1.960
Table 9: The poems generated by the CoVO-based fine-tuned model with the highest originality scores from training (top) and testing (bottom) tone-style pairs.
@@
The young man was a master of the rhyme
and in love with the young lass to this time.
So for her first night with her he wrote
a new form to try in order the note…
A true Limerick for her delight.
Oh so happy and joyous the first night of love!
$V$ = -0.818 $O$ = 1.370
@@
There was an old dog from Japan
Who fell in love with a harem of cans.
He would hide there on hot afternoons,
Till the sun turned his hair purple fawns.
One day, while watching a game in blue jeans,
He dropped by mistake the key to his knees.
He gave the other a quick whiplash and then,
Said to himself, ”Ouch! my balls are in danger!” […]
$V$ = -1.261 $O$ = 1.354
Table 10: Two limericks from the CoVO-based fine-tuned model: above, the one with the highest value score; below, the only generated one that starts with the classic first line.
Romantic haiku
spring winds
carry in the leaves.
love whispers.
$V$ = -1.343 $O$ = 2.049
Whimsical couplet
no more of one or two
let there be more.
one’s never new
but two and more’s a rarity.
$V$ = -1.361 $O$ = 1.718
Table 11: Two of the highest-scoring poems from the CoVO+KL-based fine-tuned model.